OpenStudy (freemap):

Eloisa determined that the correlation coefficient for a set of data was –0.7. What percent of the total variation in the y-values is not explained by the regression line? %

OpenStudy (freemap):

@Daniellelovee

OpenStudy (daniellelovee):

idk this one I'm sorry

OpenStudy (daniellelovee):

ask @Directrix he/she is super smart

OpenStudy (freemap):

okay thanks I didn't know he was on. I fanned you. Thanks for the help

OpenStudy (daniellelovee):

no problem and thank you :)

OpenStudy (freemap):

yw

OpenStudy (freemap):

@Directrix

OpenStudy (freemap):

@mathmale

Directrix (directrix):

•The percent of variation which is unexplained by the regression equation •The unexplained variation divided by the total variation •1 - r^2 https://people.richland.edu/james/lecture/m170/ch11-rsq.html

Directrix (directrix):

1 - r^2 = 1 - (–0.7)^2 = ? @freemap

OpenStudy (freemap):

2.4 just got the response

Directrix (directrix):

2.4 is not correct. Try again: 1 - (–0.7)^2 = 1 - (-.7) * (-.7) = ?

Directrix (directrix):

What is .7 times .7 =

OpenStudy (freemap):

.49

Directrix (directrix):

What is 1 - .49 = ?

OpenStudy (freemap):

.51

Directrix (directrix):

.51 = 51% What percent of the total variation in the y-values is not explained by the regression line

OpenStudy (freemap):

51%

Directrix (directrix):

Correct.

OpenStudy (freemap):

Thank you

Directrix (directrix):

You are welcome.